OpenStudy (sambow17):

What is the vertex for the graph of y-2=-(x+3)^2 A (-3,2) B (-2,3) C (2,-3) D (3,-2)

OpenStudy (welshfella):

the vertex form of this type of equation is y = a(x - b)^2 + c where (b,c) is the vertex

OpenStudy (welshfella):

So rearrange your equation to the above form and you'll get the answer.

OpenStudy (welshfella):

Hint: you just have to add + 2 to both sides of your equation

OpenStudy (sambow17):

So B?

OpenStudy (welshfella):

No what is the value of c when you have added the 2?

OpenStudy (sambow17):

Ha yeah its A, thanks!

OpenStudy (welshfella):

y - 2 + 2 = -(x + 3)^2 + 2

OpenStudy (welshfella):

yw

OpenStudy (welshfella):

y = a(x - b)^2 + c y = -(x + 3)^2 + 2 -b = 3 so b = -3 and c = 2

OpenStudy (welshfella):

(b , c) = (-3, 2)

OpenStudy (welshfella):

OK?